A telephone pole has three cables pulling as shown from above, with F⃗ 1=(300.0iˆ+500.0jˆ) F⃗ 2=−200.0iˆ F⃗ 3=−800.0jˆ. (a) Find the force on the telephone pole in component form. (b) Find the magnitude and direction of this net force.
1
Expert's answer
2020-01-31T10:44:25-0500

As per the given question,

Force on the pole "F_1=300.0\hat{i}+500.0\hat{j}"

"F⃗ _2=−200.0\hat{i}"

"F⃗ 3=−800.0\hat{j}"

a)

Net force on the pole "=F_1+F_2+F_3"

"=300.0\hat{i}+500.0\hat{j}−200.0\hat{i}−800.0\hat{j}=100\hat{i}-300\hat{j}"

Hence the component of the force on the pole is 100N along the x axis, and 300 along the y axis towards the ground.


b)

Net force on the pole ="|100\hat{i}-300\hat{j}|=\sqrt{100^2+300^2}=100\sqrt{10}N"

The direction of the net force "\tan\theta =\dfrac{F_y}{F_x}=\dfrac{-300}{100}=\dfrac{-3}{1}"

Hence, "\theta=71.56^\circ" downwards.